two friends started a journey for a place one by motorcycle and the other by train. the speed of the motorcycle is 30 km/h and that of train is 24 km/h ,the first take 6 h 12 min to reach the destination.find the time of reaching sond one

Grade:10

1 Answers

Utsav Basu
6 years ago
The speed of the motorcycle = 30 km/h
Time taken = 6h 12m = 6.2h
Thus, distance = 6.2 * 30km = 186km
 
The speed of train = 24 km/h
So, Time taken = (186/24)h = 7.75h = 7h 45m